a family spends 1/10 of its annual income for housing, 1/4 for food and clothing, 1/5 for general expenses, and 2/15 for entertainment. What fractional part of their income is spent on these items altogether?
Answer by stanbon(75887)

a family spends 1/10 of its annual income for housing, 1/4 for food and clothing, 1/5 for general expenses, and 2/15 for entertainment. What fractional part of their income is spent on these items altogether?
------------------------------------
(1/10)+(1/4)+(1/5)+(2/15)
lcd = 2^2*3*5 = 60
Convert each fraction:
(6/lcd) + (15/lcd) + (12/lcd) + (8/lcd)
(6 + 15 + 12 + 8) / lcd
= 41/60
==============
Cheers,
Stan H.